# It’s complicated. The relationship of algorithmic fairness and non-discrimination regulations in the EU AI Act

### v2

@@ 56 Codes of practice for GPAI Models
Furthermore, the wording of Article 10(2)(f) AIA implies an ex-post view: the developer has to know whether a bias will likely lead to discrimination. However, especially for larger models, it is technically very challenging and impossible to predict the individual output of AI models (Hacker et al., 2024b; Black et al., 2022).
@@
Article 10(2)(g) AIA: Appropriate bias detection, prevention, and mitigation on input data must include factors beyond technical means. Article 10(2)(f) AIA is completed by Article 10(2)(g) AIA. Article 10(2)(g) AIA demands that “appropriate measures to detect, prevent and mitigate possible biases identified according to point (f)” are considered. Thus, Article 10 AIA not only mandates the examination of biases that lead to discrimination but also the mitigation of these biases. The bias tests must be documented and disclosed according to Article 11 AIA, along with Annexes IV and IXa AIA (Zuiderveen Borgesius et al., 2024). It was argued that with this focus on the input side, the AI Act seeks to remedy the root cause of biases (Zuiderveen Borgesius et al., 2024).
@@
Input bias data detection, prevention, and mitigation methods remain largely unexplored. These methods are not clearly defined in the AI Act (Wachter, 2024; Deck et al., 2024b). Article 10(2)(g) AIA appears to be inspired by the technical literature on fairness metrics for algorithmic outputs. Nevertheless, Article 10(2)(g) AIA targets the input of algorithms. Therefore, classical fairness metrics for algorithmic outputs do not apply.
@@
However, compared to output fairness metrics, technical bias detection methods for input data are understudied. Furthermore, as previously noted, the causes of biases in ML are multi-causal (Gebru, 2019; Gebru et al., 2021; Hanna et al., 2020; Mehrabi et al., 2021; Bender et al., 2021). Even when data are collected representatively, biases can occur (see (Mehrabi et al., 2021) for a comprehensive list). Two examples illustrate this argument:
@@
Social bias can occur regardless of sample representativeness (Mitchell et al., 2021). For instance, in loan application settings, if certain groups (such as women) systematically receive lower wages, the data will contain inherent bias even with perfect representation.
@@
Bias can emerge when one subgroup is smaller than others (Barocas et al., 2023). Intuitively, an algorithm’s performance decreases when trained on fewer data points, resulting in poorer performance for smaller subgroups. Since smaller subgroups often correspond to minorities, algorithms may perform worse for these populations even with representative data.
@@
The data governance strategies in Article 10(2)(g) AIA focus too heavily on technical aspects of input data biases. While systematic fairness approaches have been developed for output bias detection, prevention, and mitigation (Han et al., 2023; Krasanakis and Papadopoulos, 2024; Pagano et al., 2023; Zollo et al., 2024), we lack similarly comprehensive approaches for input data analysis. This creates a significant gap between the legal requirements in Article 10(2)(g) AIA and their technical and ethical feasibility.
@@
Article 10(2) AIA: Additional obligations without specific fairness regulations In reviewing of Article 10(2) AIA, it is noticeable that the other sections (a-e) of the Article also indirectly link to algorithmic fairness considerations. For example, the representativeness of training data is explicitly mentioned in Article 10(2)(d). When interpreting Article 10(2)(f) AIA or Article 10(2)(g) AIA, however, the additional requirements in Article 10(2) AIA do not result in any new insights. Since these articles have no direct requirements regarding non-discrimination, we will not analyze them in further detail.
@@
Article 10(5) AIA: A legal basis for the processing of personal data in non-discrimination contexts. Finally, Article 10(5) AIA of the AI Act allows the processing of Article 9 GDPR data (see also Recital 70 AIA). Article 9 GDPR protects special categories of personal data, such as genetic, biometric, or health data (Article 9(1) GDPR). There is tension (Deck et al., 2024a) between the need for debiasing AI algorithms and data protection law, which Article 10(5) tries to solve. In order to effectively mitigate biases in AI systems, the processing of personal data (for example, to compute fairness metrics) is important (van Bekkum, 2024). Notable here is that this exception only applies in the high-risk regime. Thus, for non-high-risk systems, this exception is not applicable. However, it may also be a legal basis for the processing of personal data outside of high-risk systems (Hacker, 2021).
@@
Article 15 (4) AIA: Output bias regulation in case of feedback loops. In contrast to Article 10(2)(f) AIA, Article 15(4) AIA mandates that “High-risk AI systems that continue to learn after being placed on the market or put into service shall be developed in such a way as to eliminate or reduce as far as possible the risk of possibly biased outputs influencing input for future operations (feedback loops) and as to ensure that any such feedback loops are duly addressed with appropriate mitigation measures.” In this case, the output of a system is regulated, in contrast to the input of an AI system in Article 10 AIA. The reasoning behind this is that AI systems should not become echo chambers for biases (Zuiderveen Borgesius et al., 2024). According to Recital 67 AIA, this is a special concern when examining historical biases. Most important to note here is the fact, that this output regulation only applies if the systems continue learning after being placed on the market or put into service.
@@
Additionally, it remains unclear if the word “bias” here refers to the same bias as in Article 10 AIA. Recital 67 AIA, which relates to Article 10 AIA but also mentions feedback loops, indicates that the same biases are addressed.
@@
Biased output is usually detected and mitigated with fairness metrics. As shown above, the interpretation of fairness metrics is challenging. An example of this difficulty is presented in Figure 2 in the appendix. For some AI systems, there exists a trade-off between fairness and other measures such as accuracy of an AI system. Figure 2 shows a typical algorithmic fairness vs. accuracy trade-off for an AI system used to predict the income status, data are from(Mittelstadt et al., 2023). Although this trade-off can be reduced with approaches like “Levelling Up” (Mittelstadt et al., 2023), the trade-off is still present. Therefore, it becomes very challenging to identify how to “reduce as far as possible the risk of possibly biased outputs” in Article 15(4) AIA. This requirement is vague.
@@
Furthermore, especially for language (generative) models, biases in the input data and biases in the output data might be different. Furthermore, the mention of bias in Article 15, titled “Accuracy, robustness, and cybersecurity”, is rather odd since the latter two fields are traditionally not mainly concerned with discrimination (Nolte et al., 2024). Here, it remains an open question whether the robustness of an AI system and fairness represent a trade-off (Lee et al., 2021).
@@
Other articles regarding high-risk systems regulations. While the previously discussed articles provide at least some intuition about addressed concerns, we want to categorize the remaining articles of Table 1. Most of these paragraphs only concern Fundamental Rights impacts which show limited impact for non-discrimination regulation (see above). Thus, the discussion here is rather short. Article 13(3)(b)(iii) AIA mandates that information on the foreseeable misuse and Fundamental Rights impact is provided. Article 14(2) AIA mandates that human oversight is needed to minimize risks to Fundamental Rights. Article 17(1)(i) AIA and Article 26(5) AIA are related to reporting and information of serious incidents. Serious incidents are linked to discrimination through Article 3(49)(c) AIA: “’serious incident’ means an incident or malfunctioning of an AI system that directly or indirectly leads to any of the following: […] the infringement of obligations under Union law intended to protect Fundamental Rights”. The link to discrimination is even more indirect since it does not directly concern Fundamental Rights but instead Union law intended to protect Fundamental Rights. The relation to discrimination law becomes increasingly relaxed. Finally, Article 27(1) AIA requires a Fundamental Rights impact assessment when the high-risk system is used in special systems within bodies governed by public law. Yet again, significant gaps exist between theoretical and methodological elaboration of these impact assessments (Mantelero, 2024).
@@
Summary: High-risk systems balancing between specificity and generality The AI Act aims to be a specific law for AI. Classical EU non-discrimination law can already be difficult to interpret (see related literature above). In contrast to the ambitions of the AI Act, the concrete non-discrimination regulations lack specificity. For the ML community it is challenging to understand the implications of the AI Act. Only Article 10 AIA and Article 15 AIA explicitly mention non-discrimination and biases, compared to the vague statement of Fundamental Rights protection in Article 9 AIA. However, Article 10 AIA only considers the input phase, while Article 15 AIA only addresses the output phase for feedback loops. It would be beneficial if Article 15 AIA, Article 9 AIA, and Article 10 AIA regulated the same types of direct, indirect, and intersectional discrimination on both the input and output sides. In this case, the AI Act would substantially add specific benefits to classical broad-applicability non-discrimination law.
3.4. General Purpose AI non-discrimination regulations
GPAI models are an important model category. Most (Multimodal) Large Language Models fall under this category. While most of our paper focuses on high-risk regulation in the fairness domain, we briefly discuss these important GPAI models as well. Besides the high-risk systems, GPAI systems are also heavily regulated (Novelli et al., 2024).
Introduction to GPAI models. Besides high-risk systems, the AI Act regulates so-called General Purpose AI (GPAI) systems in Article 51-56 AIA  [^13] Please note that the AI Act also regulates, as a third category, certain AI systems in Article 50 AIA.. According to Article 3(6) AIA, a GPAI system is “an AI system which is based on a general-purpose AI model and which has the capability to serve a variety of purposes, both for direct use as well as for integration in other AI systems”.
GPAI models are thus models with a wide variety of capabilities that are not immediately foreseeable after training and rely on model size (Wei et al., 2022). From a technical point of view, the definition includes Large Language Models as well as Large Multimodal Models such as OpenAI’s GPT-Model family, Anthropic’s Claude models, and Google Gemini (Triguero et al., 2024). If an AI System has GPAI capabilities as well as high-risk properties, it must adhere to both GPAI and high-risk regulations (Nolte et al., 2024).
Non-discrimination regulation in GPAI models with systemic risks. GPAI models and GPAI models with systemic risk must be differentiated. The AI Act also defines GPAI models with systemic risk in Article 51(1) AIA. A GPAI model with systemic risk is either a model that has high-impact capabilities evaluated on the basis of appropriate technical tools and methodologies, including indicators and benchmarks (Article 51(1)(a) AIA), or any GPAI model whose training compute exceeds $10^{25}$ floating point operations (flops). Especially, the last condition is a topic of debate (Heim and Koessler, 2024; Novelli et al., 2024; Gupta et al., 2024). Many researchers in ML try to build models with ChatGPT-like abilities but with less compute power, for example, through quantization (Xiao et al., 2023; Gholami et al., 2022; Ma et al., 2024). Thus, the definition of GPAI model with systemic risk based on training compute is questionable. The systemic risk itself is defined in Article 3(65) AIA. A systemic risk “means a risk that is specific to the high-impact capabilities of general-purpose AI models, having a significant impact on the Union market due to their reach, or due to actual or reasonably foreseeable negative effects on public health, safety, public security, fundamental rights, or the society as a whole, that can be propagated at scale across the value chain”. The risk must be specific to the high-impact capabilities of the GPAI model. All risks that are also present in low- to high-risk systems are not covered as systemic risk. Furthermore, the impact on Fundamental Rights — and thus on non-discrimination regulation — needs to be actual or reasonably foreseeable. In common with high-risk system regulations, not all risks need to be considered but only the actual or foreseeable risks. Adding the word “reasonable” before “foreseeable” softens the non-discrimination regulation. The basic regulations for normal GPAI models are given in Article 53 AIA, while the specific regulations for GPAI models with systemic risks are stated in Article 55 AIA. However, both articles do not include any specific regulations but rather quite vague documentation, evaluation, and analysis obligations for the providers. Although there are technical measures to assess the risks of GPAI models (Paskov et al., 2024; Barrett et al., 2023; Guldimann et al., 2024), it is currently rather unclear how to deal with the GPAI regulations. The legislator has already anticipated this issue. Therefore, Article 56 AIA calls for the EU AI Office to draw up a Code of Practice together with providers and other stakeholders. Currently, this Code of Practice is being developed in four directions: Transparency and copyright-related rules, Risk identification and assessment, Technical risk mitigation, and Internal risk management and governance of General-purpose AI providers (Cod, 2024). It currently remains open whether this will result in more concrete regulations that enable technical and legal compliance.
Summary: GPAI models are only vaguely regulated. In summary, GPAI models are currently only vaguely regulated. There exists significant uncertainty, especially regarding how the regulations of the AI Act can be implemented in practice for GPAI models, besides attempts to solve this (Zollo et al., 2024). Perhaps the Code of Practice being developed by the AI Office and the European Commission can provide guidance in this area.
4. Summary & Outlook: How to proceed?
As Mayson (2018) already pointed out: Bias in, bias out. The AI act seek to complement existing non-discrimination law with it processed based regulation (Zuiderveen Borgesius et al., 2024). However, even on this broader level, the AI Act fails to account for major issues.

### v3

@@ 2 A primer on non-discrimination from an AI and a legal perspective
Furthermore, the wording of Article 10(2)(f) AIA implies an ex-post view: the developer has to know whether a bias will likely lead to discrimination. However, especially for larger models, it is technically very challenging and impossible to predict the individual output of AI models (Hacker et al. 2024; Black, Raghavan, and Barocas 2022).
@@
Article 10(2)(g) AIA: Appropriate bias detection, prevention, and mitigation on input data must include factors beyond technical means. Article 10(2)(f) AIA is complemented by Article 10(2)(g) AIA. Article 10(2)(g) AIA demands that “appropriate measures to detect, prevent and mitigate possible biases identified according to point (f)” are considered. Thus, Article 10 AIA not only mandates the examination of biases that lead to discrimination but also the mitigation of these biases. The bias tests must be documented and disclosed according to Article 11 AIA, along with Annexes IV and IXa AIA (Zuiderveen Borgesius et al. 2024). It has been argued that with this focus on the input side, the AI Act seeks to remedy the root cause of biases (Zuiderveen Borgesius et al. 2024).
@@
Article 10(2)(g) AIA appears to be inspired by the technical literature on fairness metrics for algorithmic outputs. Nevertheless, Article 10(2)(g) AIA targets the input of algorithms. Therefore, classical fairness metrics for algorithmic outputs do not apply directly. On the other hand, input bias data detection, prevention, and mitigation methods remain largely unexplored. These methods are not clearly defined in the AI Act (Wachter 2024; Deck et al. 2024b).
Even if we assume that Article 10(2)(g) AIA is applicable to the output of an algorithm, the interpretation of its meaning is not self-evident. Although the wording in Article 10(2)(g) AIA does not directly relate to fairness metrics, from a teleological interpretation of the law, one might infer that Article 10(2)(g) AIA targets traditional output-related machine learning fairness metrics. In such cases, providers must ensure that the chosen metrics are suitable for detecting possible biases. As discussed above, there are many different ways to measure fairness in machine learning systems (Mehrabi et al. 2021). Therefore, it is already difficult to select a particular method to measure fairness. Furthermore, challenges arise beyond the choice of a specific measure. Even within a single measure, trade-offs may exist between the appropriateness of fairness metrics and other requirements such as accuracy (Mittelstadt, Wachter, and Russell 2023). For example, algorithms often show an inherent trade-off between accuracy and fairness in particular settings (see Figure 1). Although this trade-off can be reduced with approaches like “Levelling Up” trade-offs are difficult to navigate. Should developers focus more on accuracy (see also Article 15 AIA) or on the non-discrimination matters? Balancing these needs is a non-trivial task. Even for computer science experts, it is currently not entirely clear how the application of fairness metrics should be managed in the context of such trade-offs. Accordingly, the question persists: What constitutes an “appropriate measure” within the meaning of the AIA? The AIA’s legislative text is sufficiently indeterminate to permit varying interpretations concerning the precise objective and intended scope of the requirement for appropriate measures.
Figure 1: Algorithmic fairness vs. accuracy trade-off, data adopted from (Mittelstadt, Wachter, and Russell 2023). While the accuracy improves, the measure of algorithmic fairness (here demographic parity for the sex attribute) increases as well. In order to reduce algorithmic unfairness, the accuracy of the algorithm decreases. Permission to replot the data was obtained from the original authors.
@@
However, compared to output fairness metrics, technical bias detection methods for input data are understudied. Furthermore, as previously noted, the causes of biases in ML are multi-causal (Gebru 2019; Gebru et al. 2021; Hanna et al. 2020; Mehrabi et al. 2021; Bender et al. 2021). Even when data are collected representatively, biases can occur (see (Mehrabi et al. 2021) for a comprehensive list). Two examples illustrate this argument:
@@
Social bias can occur regardless of sample representativeness (Mitchell et al. 2021). For instance, in loan application settings, if certain groups (such as women) systematically receive lower wages, the data will contain inherent bias even with perfect representation.
@@
Bias can emerge when one subgroup is smaller than others (Barocas, Hardt, and Narayanan 2023). Intuitively, an algorithm’s performance decreases when trained on fewer data points, resulting in poorer performance for smaller subgroups. Since smaller subgroups often correspond to minorities, algorithms may perform worse for these populations even with representative data.
@@
The data governance strategies in Article 10(2)(g) AIA focus on the technical aspects of input data biases. While systematic fairness approaches have been developed for output bias detection, prevention, and mitigation (Han, Baldwin, and Cohn 2023; Krasanakis and Papadopoulos 2024; Pagano et al. 2023; Zollo et al. 2024), we lack similarly comprehensive approaches for input data analysis. This creates a significant gap between the legal requirements in Article 10(2)(g) AIA and their technical and ethical feasibility.
@@
Article 10(2) AIA: Additional obligations without specific fairness regulations In reviewing of Article 10(2) AIA, it is noticeable that the other sections of Article 10(2) AIA also indirectly link to algorithmic fairness considerations. For example, the representativeness of training data is explicitly mentioned in Article 10(2)(d). When interpreting Article 10(2)(f) AIA or Article 10(2)(g) AIA, however, the additional requirements in Article 10(2) AIA do not result in any new insights. Since these articles have no direct requirements regarding non-discrimination, we will not analyze them in further detail.
@@
Article 10(5) AIA: A legal basis for the processing of personal data in non-discrimination contexts. Finally, Article 10(5) AIA allows the processing of Article 9 GDPR data (see also Recital 70 AIA). Article 9 GDPR protects special categories of personal data, such as genetic, biometric, or health data (Article 9(1) GDPR). There is tension (Deck et al. 2024a) between the need for debiasing AI algorithms and data protection law, which Article 10(5) aims to address. In order to effectively mitigate biases in AI systems, the processing of personal data (for example, to compute fairness metrics) is important (van Bekkum 2024). Notable here is that this exception only applies in the high-risk regime. Thus, this exception is not applicable to non-high-risk systems. Nevertheless, Article 10(5) is an example of how the AI Act not only imposes burdens on developers and deployers of AI systems but also grants rights to these groups. This aspect is frequently underappreciated in the discourse regarding the AI Act.
Article 11 (1) AIA and Annex IV (2)(g) AIA: Technical documentation of bias testing methods
Article 11 (1) together with Annex IV (2)(g) AIA requires that the technical documentation of a high-risk system includes “information about the validation and testing data used and their main characteristics; metrics used to measure accuracy, robustness and compliance with other relevant requirements set out in Chapter III, Section 2, as well as potentially discriminatory impacts”.
From a grammatical interpretation, this section only specifies that the methods enforced in Chapter III, Section 2 AIA are documented. No new additional requirements are enforced. Would a teleological interpretation change this analysis? Even if Article 11(1) AIA, together with Annex IV (2)(g) AIA, were to enforce new measures, the requirement would again only target validation and testing data  [^13] In contrast to Article 10, AIA training data are not mentioned. . Thus, no additional and new requirements for high-risk AI systems on the output side are enforced through Article 11(1) AIA.
Article 13 AIA: Transparency and Provision of Information to Deployers Article 13(1) AIA provides that ‘High-risk AI systems shall be designed and developed in such a way as to ensure that their operation is sufficiently transparent to enable deployers to interpret a system’s output and use it appropriately”. One could contend that the interpretation of a system’s output necessarily entails an assessment of potential discrimination. Recital 72 enumerates aspects regarding the information that should be provided, with specific reference to the protection of fundamental rights. However, the risk identified in the recital is presented in general terms and fails to specify any particular methodology or technological tool that must be employed. Moreover, as Recitals serve a non-binding role and do not possess direct normative force in the interpretation of Article 13, the lack of specificity in Recital 72 does not alter the conclusion. Accordingly, Article 13(1) itself does not impose an obligation to use algorithmic fairness measures or metrics derived from the computer science domain.
Furthermore, even assuming, arguendo, that the use of such algorithmic fairness tools is required during the development or deployment process, Article 13(1) merely obliges providers to document the results. The imposition of mitigation or prevention measures does not fall within the mandatory scope of Article 13(1).
@@
Article 15 (4) AIA: Output bias regulation in case of feedback loops. In contrast to Article 10(2)(f) AIA, Article 15(4) AIA mandates that “High-risk AI systems that continue to learn after being placed on the market or put into service shall be developed in such a way as to eliminate or reduce as far as possible the risk of possibly biased outputs influencing input for future operations (feedback loops) and as to ensure that any such feedback loops are duly addressed with appropriate mitigation measures.” In this case, the output of a system is regulated, in contrast to the input of an AI system in Article 10 AIA. The reasoning behind this is that AI systems should not become echo chambers for biases (Zuiderveen Borgesius et al. 2024). According to Recital 67 AIA, this is a special concern when examining historical biases. Most important to note here is the fact, that this output regulation only applies if the systems continue learning after being placed on the market or put into service.
@@
Additionally, it remains unclear if the word “bias” refers to the same bias as in Article 10 AIA. Recital 67 AIA, which relates to Article 10 AIA but also mentions feedback loops, indicates that the same biases are addressed.
@@
Biased output is usually detected and mitigated with fairness metrics. As shown above, the interpretation of fairness metrics is challenging as discussed above. Therefore, it becomes very challenging to identify how to “reduce as far as possible the risk of possibly biased outputs” in Article 15(4) AIA. This requirement is vague.
@@
Furthermore, especially for language (generative) models, biases in the input data and biases in the output data might be different. In this context, the mention of bias in Article 15, titled “Accuracy, robustness, and cybersecurity”, is rather odd since the latter two fields are traditionally not mainly concerned with discrimination (Nolte, Rateike, and Finck 2024). Here, it remains an open question whether the robustness of an AI system and fairness represent a trade-off (Lee et al. 2021).
@@
Other articles regarding high-risk systems regulations. While the previously discussed articles provide at least some intuition about addressed concerns, we categorize the remaining articles of Table 1. Most of these paragraphs only concern fundamental rights impacts, which show a limited impact on non-discrimination regulations. Thus, the discussion here is rather short. Article 13(3)(b)(iii) AIA mandates that information on the foreseeable misuse and the impact on fundamental rights is provided. Article 14(2) AIA mandates that human oversight is needed to minimize risks to fundamental rights. Article 17(1)(i) AIA and Article 26(5) AIA are related to reporting and information on serious incidents. Serious incidents are linked to discrimination through Article 3(49)(c) AIA: “’serious incident’ means an incident or malfunctioning of an AI system that directly or indirectly leads to any of the following: […] the infringement of obligations under Union law intended to protect fundamental rights”. Finally, Article 27(1) AIA requires a fundamental rights impact assessment when the high-risk system is used in special systems within bodies governed by public law. Yet again, significant gaps exist between theoretical and methodological elaboration of these impact assessments (Mantelero 2024). Article 27(1) AIA only addresses public parties. A fundamental right impact assessment is not necessary for private parties.
@@
Finally, the impact of fundamental rights on non-discrimination is also monitored by a specific authority established through Article 77 (1) and Article 77 (2) AIA. The effect and impact of this authority framework should be assessed and evaluated after it is established. Summary: High-risk systems balancing between specificity and generality The AI Act aims to be a specific law for AI (Ződi 2024). For the ML and Law community, it might be challenging to understand the implications of this specific AI Law. Some Articles mention non-discrimination directly, while others focus on fundamental rights. Only Article 10 AIA and Article 15 AIA explicitly mention non-discrimination and biases, compared to the statement of fundamental rights protection in Article 9 AIA. However, Article 10 AIA only considers the input phase, while Article 15 AIA only addresses the output phase for feedback loops. It remains open why different Articles in the AI Act do not regulate the same types of direct, indirect, and intersectional discrimination on both the input and output sides.
Interplay between traditional EU non-discrimination law and the AI Act for high-risk systems
The AI Act focuses on the input side of high-risk systems. The output side is not covered in comparable detail. From a computer science perspective, research focuses on the output side of algorithms. As a lex specialis (Craig and De Búrca 2011), the AI Act prevails over general frameworks to the extent that it covers an issue; where it does not address certain discriminatory outcomes, however, the more general non-discrimination law remains applicable. Since output regulation is not present, traditional EU non-discrimination law could be applied.
The interplay between the AI regulations and fundamental rights requires further research, including clarification regarding regulation when the AI Act protects specific fundamental rights such as non-discrimination (Lewis et al. 2025). This also includes an analysis of when fundamental rights are vertically and horizontally applicable. Furthermore, the fundamental rights in the AI Act need to be balanced (Kusche 2024), and the material scope is limited due to the enumeration of protected attributes (Xenidis and Senden 2020). As pointed out, “gaps in the regulatory framework have left fundamental rights inconsistently protected in the AI Act.” (Palmiotto 2025) Nevertheless, since the AI Act does not protect individuals regarding the output of algorithms, there is room for the applicability of the CFR. It remains an open research question how the protection in the CFR and the AI Act relate (Lewis et al. 2025).
Besides the CFR, the protection of individuals from non-discrimination has been concretized in secondary EU law (Xenidis and Senden 2020; Lewis et al. 2025). Compared to the CFR, the horizontal applicability is clear. For example, Directives 2004/113/EC (Equal Treatment in Goods and Services Directive) and 2000/43/EC (Race and Ethnicity Equality Directive) specifically target non-discrimination regulations. These regulations address specific cases in which discrimination can occur. Due to their specificity, they have a limited material and personal scope as a result of the enumeration of protected attributes (Xenidis and Senden 2020; Cîrciumaru 2024). . It therefore remains to be seen how the interplay between the existing regulations, which cannot simply be transferred to AI, and the AI Act will develop. It also remains to be seen how national and EU courts will position themselves. In the past, the ECJ has already issued important rulings on how the non-discrimination provisions in Article 21 CFR must be interpreted in relation to secondary EU law (Xenidis and Senden 2020). It is very likely that in the future, the ECJ will continue to issue rulings to clarify the interplay between the EU AI Act, fundamental rights, and other directives and regulations. There is still a long way to go.
The AI Act holds the potential to complement traditional non-discrimination legislation. Traditional non-discrimination law primarily targets the protection of individuals. Despite the identified challenges associated with this individual-focused approach, it nonetheless provides a degree of protection. However, such an emphasis on individuals can overlook broader regulatory dimensions. In the context of EU data protection, it has been posited that a dual approach involving both individual protection and collaborative governance is necessary (Kaminski 2018). A dual approach might also be beneficial for non-discrimination regulations in the context of AI. The EU AI Act is fundamentally aligned with product safety regulation rather than individual rights protection (Almada and Petit 2023). The strength of this regulatory framework lies in its capacity to extend beyond individual-focused concerns. To maximize its efficacy, we contend that it is imperative to clarify the relationship between this framework and traditional non-discrimination regulations. It is plausible that both the product safety focus and individual rights considerations will be key requirements for the effective and comprehensive regulation of AI systems.
4 Summary & Outlook
As Mayson (2018) already pointed out: Bias in, bias out. While regulating these biases, AI encounters significant complexities.